OpenStudy (anonymous):

Solve the following: (Show work please) |x - 5| = 4 |x - 5| < 4

OpenStudy (anonymous):

what have you tried ?

OpenStudy (anonymous):

For the first one, I've done x - 5 + 5 = 4 + 5 So x = 9, but that's not one of the options for answers.

OpenStudy (anonymous):

I'm supposed to match it with it's "equivalent compound sentence"

OpenStudy (anonymous):

so i'm not exactly sure what i'm doing lol

OpenStudy (anonymous):

These are simultaneous or single equations?

OpenStudy (anonymous):

Match each absolute value sentence with its equivalent compound sentence. Match |x - 5| = 4 |x - 5| < 4 |x - 5| > 4 |2x + 3| < 8 |2x + 3| > 8 With A. x - 5 > 4 or x - 5 < -4 B. -4 < x - 5 < 4 C. -8 < 2x + 3 < 8 D. x - 5 = 4 or x - 5 = -4 E. 2x + 3 > 8 or 2x + 3 < -8

OpenStudy (anonymous):

cool they are single, so... what do we know about |x|< a ?

OpenStudy (anonymous):

Umm, x is less than a? (Bare with me I'm horrible at math lol)

OpenStudy (anonymous):

\( |x-b| < a \Rightarrow -a<x-b<a \) so, \(|x - 5| < 4 \Rightarrow 1<x<9 \)

OpenStudy (anonymous):

Hence, |x - 5| < 4 is B

OpenStudy (anonymous):

I got it now thanks